OpenStudy (anonymous):

10 algebraic expressions?

OpenStudy (anonymous):

99 bottles of beer on the wall ?

OpenStudy (anonymous):

gimme c:

OpenStudy (anonymous):

What are your expressions?

OpenStudy (anonymous):

\(2x\) \(3-y\) \(5x - 8y\) \[\frac{x}{y} -\frac{1}{x}\] \((xy)^4\) \[4^x\] They don't have to have variables either \(2\) \(5^{1/2}\) \[\sqrt[3]{27}\] \(2 \cdot 5\) \[\frac{5}{9}\]

OpenStudy (anonymous):

thank you! :D